# Estimating the detection of antineutrinos in the future Brazilian neutron source

### v2

@@ Abstract
In line with the tradition of employing nuclear reactors for neutrino physics research, we present potential experiments that could be undertaken using the new Brazilian neutron source, the Brazilian Multipurpose Reactor (RMB). This upcoming facility has clearly defined objectives, including the production of radiopharmaceuticals for the healthcare system, material irradiation testing, and research using neutron beams to investigate the micro and mesostructures of various materials. However, we anticipate that the RMB will generate up to $3,500$ antineutrinos daily, potentially expanding the facility’s scope to include neutrino physics experiments involving beta and inverse beta reactions. The modern operational conditions of the RMB and its location in the Iperó-Brazil region may also enable neutrino oscillation experiments and the determination of the $\theta_{13}$ parameter, as well as support safeguards activities by the International Atomic Energy Agency (IAEA), allowing for the monitoring of nuclear fuel burning cycles and isotope fractions. We discuss the potential utilization of the RMB as an antineutrino source, proposing experiments and addressing associated technical limitations.
@@ 1 Introduction
The attribution of a finite mass to neutrinos serves as a compelling indication of physics beyond the standard model of elementary particles [6], which currently undergoes numerous theoretical extensions, including supersymmetry [7], novel models for dark matter [8], and superstrings [9], among others [10, 11]. Neutrino oscillation, a phenomenon of significant interest within the scientific community [12, 13], led to the independent discoveries by physicists Arthur McDonald and Takaaki Kajita, meriting them the 2015 Nobel Prize in Physics. Despite this progress, several fundamental properties of neutrinos remain unanswered, including the mass hierarchy among the three flavors and the $\theta_{13}$ parameter, crucial for describing flavor oscillations [13, 5]. Presently, numerous large-scale international collaborations focus on neutrino experiments, such as NOvA [14], MINOS [15], and DUNE [16] at Fermilab-USA, T2K [17] and Super-Kamiokande [18] in Japan, Double Chooz [19] and RICOCHET [20] in France, and SNO [21] in Canada. In Brazil, two notable neutrino experiments are conducted in the city of Angra dos Reis, Rio de Janeiro state, namely the CONNIE collaboration (Coherent Neutrino-Nucleus Interaction Experiment) [22] and ANGRA [23].
@@
The CONNIE and ANGRA experiments represent unique endeavors in the global Southern Hemisphere, sharing a common neutrino source, the $4$ $GW$ Angra2 nuclear reactor. CONNIE is dedicated to investigating coherent neutrino-nucleus interactions to unveil physics beyond the standard model, employing CCD detectors (Charge-Coupled Devices), sophisticated instruments capable of manipulating charges in small pulses [24]. Initially focused on studying neutrino oscillation, ANGRA employed traditional tank-based techniques with ultrafiltered water, surrounded by photomultipliers [25, 26, 27]. However, due to budget constraints, the collaboration shifted its focus to reactor core monitoring, utilizing a domestically produced detector. Leveraging the linear correlation between a nuclear reactor’s thermal power and neutrino production [28], the experiment facilitates real-time monitoring of Angra2, thereby contributing to the safeguard mission of the International Atomic Energy Agency (IAEA) [29].
@@
Brazil is on the verge of constructing its new neutron source, the Brazilian Multipurpose Reactor (RMB) [30]. Situated in the city of Iperó, within the interior of São Paulo State, this research reactor will cater to the entire Brazilian populace, facilitating the production of radiopharmaceuticals for both the Unified Health System (SUS) and the private sector, alongside conducting irradiation tests on materials and engaging in neutron beam studies across a diverse array of techniques and materials [31]. The fuel element powering the RMB is domestically developed through the scientific endeavors of the Institute of Energy and Nuclear Research (IPEN) in São Paulo city, where seasoned researchers have formulated methodologies for constructing and analyzing the device [32], with a thermal power output of $30$ $MW$. Functioning as the dedicated facility for neutron-based instruments, the National Neutron Laboratory (LNN) will utilize thermal and cold neutron beams as probes to investigate the mesoscopic structure of matter, offering auxiliary laboratories for sample preparation and analysis, accessible to both industry and the broader Brazilian scientific community [31]. The LNN is poised to host a suite of $15$ instruments, including Small-Angle Neutron Scattering (SANS), featuring High-Intensity [33] and High-Resolution Diffractometers, in addition to a Neutron Imaging Center [34], among other facilities [31]. Currently, the project has advanced to the stage of finalizing detailed engineering designs and initiating groundwork on the designated site.
@@ 2.1 Theoretical Aspects
$U_{MNS}=\begin{pmatrix}1&0&0\\ 0&c_{23}&s_{23}\\ 0&-s_{23}&c_{23}\end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{i\delta}\\ 0&1&0\\ -s_{13}e^{-i\delta}&0&c_{13}\end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\end{pmatrix},$ (2.2)

### v3

@@ Abstract
In line with the tradition of employing nuclear reactors for neutrino physics research, we present potential experiments that could be undertaken using the new Brazilian neutron source, the Brazilian Multipurpose Reactor (RMB). This upcoming facility has clearly defined objectives, including the production of radioisotopes for the healthcare system, material irradiation testing, and research using neutron beams to investigate the micro and mesostructures of various materials. However, we anticipate that the RMB will generate up to $3,500$ antineutrinos daily, potentially expanding the facility’s scope to include neutrino physics experiments involving beta and inverse beta reactions. The modern operational conditions of the RMB and its location in the Iperó-Brazil region may also enable neutrino oscillation experiments and the determination of the $\theta_{13}$ parameter, as well as support safeguards activities by the International Atomic Energy Agency (IAEA), allowing for the monitoring of nuclear fuel burning cycles and isotope fractions. We discuss the potential utilization of the RMB as an antineutrino source, proposing experiments and addressing associated technical limitations.
@@ 1 Introduction
The attribution of a finite mass to neutrinos serves as a compelling indication of physics beyond the standard model of elementary particles [6], which currently undergoes numerous theoretical extensions, including supersymmetry [7], novel models for dark matter [8], and superstrings [9], among others [10, 11]. Neutrino oscillation, a phenomenon of significant interest within the scientific community [12, 13], led to the independent discoveries by physicists Arthur McDonald and Takaaki Kajita, meriting them the 2015 Nobel Prize in Physics. Despite this progress, several fundamental properties of neutrinos remain unanswered, including the mass hierarchy among the three flavors and the $\theta_{13}$ parameter, crucial for describing flavor oscillations [13, 5]. Presently, numerous large-scale international collaborations focus on neutrino experiments, such as NOvA [14], MINOS [15], and DUNE [16] at Fermilab-USA, T2K [17] and Super-Kamiokande [18] in Japan, Double Chooz [19] and RICOCHET [20] in France, SNO [21] in Canada, and IceCube in South Pole [22]. In Brazil, two notable neutrino experiments are conducted in the city of Angra dos Reis, Rio de Janeiro state, namely the CONNIE collaboration (Coherent Neutrino-Nucleus Interaction Experiment) [23] and ANGRA [24].
@@
The CONNIE and ANGRA experiments are the only endeavors that use power reactor (the $4$ $GW$ Angra2) as neutrino source in the global Southern Hemisphere. CONNIE is dedicated to investigating coherent neutrino-nucleus interactions to unveil physics beyond the standard model, employing CCD detectors (Charge-Coupled Devices), sophisticated instruments capable of manipulating charges in small pulses [25]. Initially focused on studying neutrino oscillation, ANGRA employed traditional tank-based techniques with ultrafiltered water, surrounded by photomultipliers [26, 27, 28]. However, due to budget constraints, the collaboration shifted its focus to reactor core monitoring, utilizing a domestically produced detector. Leveraging the linear correlation between a nuclear reactor’s thermal power and neutrino production [29], the experiment facilitates real-time monitoring of Angra2, thereby contributing to the safeguard mission of the International Atomic Energy Agency (IAEA) [30].
@@
Brazil is on the verge of constructing its new neutron source, the Brazilian Multipurpose Reactor (RMB) [31]. Situated in the city of Iperó, within the interior of São Paulo State, this research reactor will cater to the entire Brazilian populace, facilitating the production of radioisotopes for both the Unified Health System (SUS) and the private sector, alongside conducting irradiation tests on materials and engaging in neutron beam studies across a diverse array of techniques and materials [32]. The fuel element powering the RMB is domestically developed through the scientific endeavors of the Institute of Energy and Nuclear Research (IPEN) in São Paulo city, where seasoned researchers have formulated methodologies for constructing and analyzing the device [33], with a thermal power output of $30$ $MW$. Functioning as the dedicated facility for neutron-based instruments, the National Neutron Laboratory (LNN) will utilize thermal and cold neutron beams as probes to investigate the mesoscopic structure of matter, offering auxiliary laboratories for sample preparation and analysis, accessible to both industry and the broader Brazilian scientific community [32]. The LNN is poised to host a suite of $15$ instruments, including Small-Angle Neutron Scattering (SANS), featuring High-Intensity and High-Resolution Diffractometers [34, 35], in addition to a Neutron Imaging Center [36], among other facilities [32]. Currently, the project has advanced to the stage of finalizing detailed engineering designs and initiating groundwork on the designated site.
@@ 2.1 Theoretical Aspects
$U_{MNS}=\begin{pmatrix}1&0&0\\ 0&c_{23}&s_{23}\\ 0&-s_{23}&c_{23}\end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{i\delta}\\ 0&1&0\\ -s_{13}e^{-i\delta}&0&c_{13}\end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\end{pmatrix}$ (2.2)